If you are in the trial and error stage of a ROR analysis and got the following answer: Let Ai* = 11% -15,000 + 5,000(P/A,
Answer #1

Ans part c) 8%

According to questions,

-15000 + 5000(P/A, 11%, 4) = 512

=) 512+15000 =5000 (P×11×4)/100

=) 15512/50=P×44

=) 7.05

=) 8 approximately